# ViM-Disparity: Bridging the Gap of Speed, Accuracy and Memory for Disparity Map Generation

### v1

@@ I Introduction
  • A robust benchmarking of the recent DMG techniques available in the literature. .
@@ II Proposed Methodology
After a critical study of emerging works in real-time disparity map computation, we observed that an attention-based model is quite suitable for such application [15]. Joint attention between extracted features from both left and right images, rather than the raw image itself has been explored in such models. Though attention models are quite heavy and thus take a considerable time to compute the disparity map for a pair of stereo frames, their accuracy is very robust. The recent introduction of Mamba [36], a direct state space model (SSM) an alternative for transformers that can help in reducing the computation time for DMG. Due to the linear nature of state scale compared to the quadratic nature in transformers, the overall memory footprint and computational load of Mamba are significantly less for higher input sizes. Our proposed methodology introduces to use of the advantage of both the attention mechanism and SSM, by replacing that transformer block with a Vision Mamba (ViM) [35] unit for DMG.
@@ II-A Proposed Architecture
where $\mathcal{V}_{\text{block}}$ represents the transformation applied by the ViM block at the $l$-th layer.
After processing through the ViM block, the output sequences $\tilde{F}_{\text{output}}$ and $\tilde{F}_{\text{rev, output}}$ are split to separate the left and right feature sequences. Since the concatenated features are symmetrical, we split them evenly:
@@
where $N$ is the length of the feature sequence for each image.
Global and local feature matching is then performed on the separated left and right feature sequences $F_{\text{left, output}}$ and $F_{\text{right, output}}$ to estimate the disparity map. The disparity map is further refined using regressive refinement and upsampling, which applies convex sampling to smooth the final output. Finally, we retain only the positive disparities and upsample the predicted disparity map to the original image resolution.
II-B Proposed Performance Measure
@@ IV Experiments and Analysis
TABLE I: Comparative performance of different metrics over different datasets.
@@
TABLE III: Table showing the ablation study
@@ V Conclusion
Disparity map generation (DMG) is a well-explored area in the robotic vision community. The literature on DMG exhibits that high accuracy and real-time generation with low computation overhead are found to be a trade-off. Further, most existing techniques aim to enhance the accuracy or speed of DMG. In contrast, this work attempted to extensively analyse the aforementioned trade-off of real timeliness vs accuracy of recent DMG methodologies on standard datasets such as KITTI, Virtual KITTI, Sceneflow and Sintel. Concluding from this we propose a ViM-based model which can bridge the gap of time, memory requirement and accuracy of a real-time DMG.
Moreover, we introduce a new measure for the overall joint analysis of the efficacy of DGM models by considering accuracy, speed and memory footprint. The results and analysis conclude that we were able to dissolve the gap in speed, accuracy and memory better than any other state-of-the-art techniques as per our proposed measure for DMG.

### v2

@@ I Introduction
  • A robust benchmarking of the recent DMG techniques available in the literature.
@@ II Proposed Methodology
After a critical study of emerging works in real-time disparity map computation, we observed that an attention-based model is quite suitable for such application [15]. Joint attention between extracted features from both left and right images, rather than the raw image itself has been explored in such models. Though attention models are quite heavy and thus take a considerable time to compute the disparity map for a pair of stereo frames, their accuracy is very robust. The recent introduction of Mamba [36], a direct state space model (SSM) an alternative for transformers that can help in reducing the computation time for DMG. Due to the linear nature of state scale compared to the quadratic nature in transformers, the overall memory footprint and computational load of Mamba are significantly less for higher input sizes. Our proposed methodology introduces the use of the advantage of both the attention mechanism and SSM, by replacing that transformer block with a Vision Mamba (ViM) [35] unit for DMG.
@@ II-A Proposed Architecture
where $\mathcal{V}_{\text{block}}$ represents the transformation applied by the ViM block at the $l$-th layer. After processing through the ViM block, the output sequences $\tilde{F}_{\text{output}}$ and $\tilde{F}_{\text{rev, output}}$ are split to separate the left and right feature sequences. Since the concatenated features are symmetrical, we split them evenly:
@@
where $N$ is the length of the feature sequence for each image. Global and local feature matching is then performed on the separated left and right feature sequences $F_{\text{left, output}}$ and $F_{\text{right, output}}$ to estimate the disparity map. The disparity map is further refined using regressive refinement and upsampling, which applies convex sampling to smooth the final output. Finally, we retain only the positive disparities and upsample the predicted disparity map to the original image resolution.
II-B Matching
For depth estimation, with known camera parameters (K1, E1, K2, E2) for images I1 and I2, we discretize a predefined depth range into N candidates.For each depth candidate $d_{i}$, we compute 2D correspondences in $Frame2_{2}(F_{2})$ using the transformation H($\hat{G}_{2D}$) = K2E2E1-1diK1-1H($G_{2D}$), then perform bilinear sampling to obtain feature correlations $C=F1\textperiodcentered Fi2/\sqrt{D}$.These correlations are concatenated into $C_{d}$, normalized via softmax to get the matching distribution $M_{d}$, and the final depth is estimated as $V_{d}$ by computing the weighted average of $M_{d}$ with the depth candidates $G_{d}$.
II-C Proposed Performance Measure
@@ IV Experiments and Analysis
TABLE I: Comparative performance over different datasets.
@@
TABLE III: Ablation study on different component of proposed model.
@@ V Conclusion
Disparity map generation (DMG) is a well-explored area in the robotic vision community. The literature on DMG exhibits that high accuracy and real-time generation with low computation overhead are found to be a trade-off. Further, most existing techniques aim to enhance the accuracy or speed of DMG. In contrast, this work attempted to extensively analyse the aforementioned trade-off of real timeliness vs accuracy of recent DMG methodologies on standard datasets such as KITTI, Virtual KITTI, Sceneflow and Sintel. Concluding from this we propose a ViM-based model which can bridge the gap of time, memory requirement and accuracy of a real-time DMG. Moreover, we introduce a new measure for the overall joint analysis of the efficacy of DGM models by considering accuracy, speed and memory footprint. The results and analysis conclude that we were able to dissolve the gap in speed, accuracy and memory better than any other state-of-the-art techniques as per our proposed measure for DMG.
Acknowledgment
This work was partially funded by BITS Pilani New Faculty Seed Grant (NFSG/HYD/2023/H0831).